OpenStudy (liizzyliizz):

if f(x)= x^3-x^2+x show that there is a number c such that f(c)=10

OpenStudy (anonymous):

really? oh i bet this is a "intermediate value" theorem problem isn't it?

OpenStudy (liizzyliizz):

the one after this is, and im reading on it as we speak.. im just a bit lost

OpenStudy (liizzyliizz):

i guess this is one as well, but it doesnt specify.

OpenStudy (liizzyliizz):

is*

OpenStudy (anonymous):

really not much to show because \[f(x)=x^3-x^2+x\] is continuous and has range all real numbers. so not only does it equal 10 somewhere but it equal every number somewhere

OpenStudy (anonymous):

if you want to be a little more specific (which is silly) you could say that \[f(0)=0, f(3)=21\] and since \[0<10<21\] there must be a number c in the interval (0,3) for which \[f(x)=10\]

OpenStudy (anonymous):

"intermediate value theorem" basically says that a continuous function cannot skip values. so if \[f(x_1)=a\] and \[f(x_2)=b\] then for any number with c \[a<c<b\] there is an \[x_3\] with \[f(x_3)=c

OpenStudy (anonymous):

of course i am assuming that \[a<b\] othewise flip the inequality

OpenStudy (liizzyliizz):

you just helped me on my next problem lol but thats it ? just plug in a number for example is has (0,1) for the function is x^3-3x+1

OpenStudy (anonymous):

not sure what the question is. are you given a specific value?

OpenStudy (liizzyliizz):

it says use the intermediate value theroem to show that there is a root of the given equation in the specific interval. x^3-3x+1=0 , (0,1)

OpenStudy (anonymous):

right

OpenStudy (liizzyliizz):

the whole root thing kind of threw me off...

OpenStudy (anonymous):

plug in 0 for x, see that you get some negative number

OpenStudy (anonymous):

plug in 1 for x, see that you get some positive number

OpenStudy (anonymous):

and since 0 is between whatever negative number you get and whatever positive number you get, you can conclude that there is some c in (0,1) with f(c)=0

OpenStudy (anonymous):

"root" in this case means a solution to \[x^3-3x+1=0\]

OpenStudy (anonymous):

clear or no?

OpenStudy (liizzyliizz):

yes, its clear, so i did and for 0 i got 1 and for 1 i got -1. so c would be between -1 and 1?